OpenStudy (anonymous):

How many solutions does the system of equations have? x-3y=15 and y=1/3x-5 A.one B.two C.non D. Infinitely many Please help me figure this out correctly.

OpenStudy (campbell_st):

write the 1st line in slope intercept form what do you notice..?

OpenStudy (anonymous):

I'm not sure. I think I am doing it incorrectly, I did something like 3y=15-x

OpenStudy (campbell_st):

well that's a start but its -3y = 15 - x now divide every term by -3 what do you get

OpenStudy (anonymous):

I got y=1/3x-5

OpenStudy (campbell_st):

ok... and what do you notice now about the 2 lines...?

OpenStudy (anonymous):

I notice one line.... They are on top of eachother.

OpenStudy (anonymous):

So that would be two, correct??

OpenStudy (campbell_st):

thats correct... so every point on the 1st line is on the 2nd line... so there are an infinite number of solutions.

OpenStudy (anonymous):

Thanks campbell_st for taking time out of your day to help me. I am terrible at math as you can now tell.

OpenStudy (campbell_st):

well its just practice... good luck

OpenStudy (anonymous):

Definitely practice. I need to keep looking at these types of questions. I have tons more so, it should work out fine.